OpenStudy (anonymous):

antidifferentiate : 2xe^(-x^2)

OpenStudy (anonymous):

\[\int\limits_{}^{}2xe ^{-x ^{2}}\]

OpenStudy (john_es):

\[-e^{x^2}\]

OpenStudy (anonymous):

how do u do it?

OpenStudy (john_es):

Well, it is an inmediate integral, because \[\int g'(x)f(g(x))dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}2xe ^{-x ^{2}} = -\int e^{-x^2}d(-x^2)\]

OpenStudy (john_es):

But you can try with a change of variable, \[x^2=t\\2xdx=dt\]

OpenStudy (anonymous):

so what happens if its g'x f(g(x)?

OpenStudy (john_es):

In this particular case, we have \[\int g'(x)f'(g(x))dx=\int \frac{d(f(g(x)))}{dx} dx=f(g(x))\]

OpenStudy (john_es):

When you have integrals that cannot solve this way, but still are "inmediate", you can do a change of variable. I suggested x^2=t, but watching the answer of @wio you can also do, -x^2=t. \[-x^2=t\\ -2xdx=dt\Rightarrow2xdx=-dt\\ \int 2xe^{x^2}dx=-\int e^tdt=-e^t=-e^{x^2}\]Don't forget to add the usual constant for indefinite integrals, \[-e^{x^2}+K\]

OpenStudy (anonymous):

Thankyou

OpenStudy (john_es):

;)